# Examining Multilingual Embedding Models Cross-Lingually Through LLM-Generated Adversarial Examples

### v1

@@ Abstract
The evaluation of cross-lingual semantic search capabilities of models is often limited to existing datasets from tasks such as information retrieval and semantic textual similarity. To allow for domain-specific evaluation, we introduce Cross Lingual Semantic Discrimination (CLSD), a novel cross-lingual semantic search task that requires only a set of parallel sentence pairs of the language pair of interest within the target domain. This task focuses on the ability of a model to cross-lingually rank the true parallel sentence higher than hard negatives generated by a large language model. We create four instances of our introduced CLSD task for the language pair German-French within the domain of news. Within this case study, we find that models that are also fine-tuned for retrieval tasks (e.g., multilingual E5) benefit from using English as the pivot language, while bitext mining models such as LaBSE perform best directly cross-lingually. We also show a fine-grained similarity analysis enabled by our distractor generation strategy, indicating that different embedding models are sensitive to different types of perturbations.
@@ 1 Introduction
When considering which model to use to perform semantic search within a specific domain and language pair, standard benchmarks such as MTEB Muennighoff et al. (2023) do not cover every domain within all language pairs. We posit that one should aim to directly evaluate semantic search for the target language pair within the domain of interest but the lack of dataset for such evaluation cases makes it difficult to take fundamental decisions for the application uses, such as the choice of an embedding model or whether to use English as a pivot language to perform the task.
@@ 2 Related Work
PAWS-X Yang et al. (2019); Zhang et al. (2019) generated adversarial examples by rule-based and backtranslation methods to measure the limitations of paraphrase identification models. More recently, XSim++ Chen et al. (2023) used rule-based text augmentation to generate synthetic examples for tuning LASER models (replacements by antonyms, swapping entities). Within semantic search, InPars (Bonifacio et al., 2022) use solely in-domain synthetic data to finetune retrieval models that surpass strong baselines.
@@
| Machine Translation (M2M 1.2B Fan et al. (2021)) to EN -> CLSD Evaluation | ||||||||
@@ 3.1 Cross-Lingual Semantic Discrimination (CLSD)
In this paper, we test the hypothesis that LLMs can generate difficult examples that look similar on the surface, but convey different meanings. We prompt GPT-4 (gpt-4-0613) OpenAI (2023) to monolingually generate four distractors for each original sentence (see Figure 1). The distractors are requested to be structurally and lexically similar, whilst being semantically dissimilar. For each language, we manually examine 200 distractors of the original sentences and find that over 98% of the distractors meet these criteria. Example sentences are shown in Table 1.
@@
We generate and validate (samples of) four datasets, consisting of both language directions of German/French for the WMT19 Barrault et al. (2019) and WMT21 Akhbardeh et al. (2021) test sets.111The data is available under ANONYMURL.. In the 18,928 generated distractors, we observe that the LLM used different modification strategies such as single polarity inversion, entity replacement or even multiple swaps with word-level Jaccard similarity between originals and distractors at (, ). Table 4 illustrates samples of distractors sets.
@@
To measure performance of embedding models, we use Recall@1 (R@1), defined as the proportion of samples in which the model ranks the true translation as more similar to the source sentence than the four distractors. For example, in Table 2, the column WMT19-DE->FR reports the R@1 of identifying the original French target sentence as more similar than its French distractors.
@@
We examine the following five embedding models available through SentenceTransformers Reimers and Gurevych (2019).
multilingual-e5-base/-large (M-E5-b/l) Wang et al. (2024, 2024) transform XLM-RoBERTa Conneau et al. (2020) into a bi-encoder through weakly supervised contrastive pre-training with multilingual data and a second step of supervised fine-tuning on retrieval datasets.
@@
paraphrase-multilingual-mpnet-base-v2 (M-MPNet) Reimers and Gurevych (2020) use knowledge distillation through cosine loss training on parallel sentences between an English paraphrase trained teacher model (paraphrase-mpnet-base-v2) and a multilingual student model (XLM-RoBERTa Conneau et al. (2020)) to teach the student to represent sentences closely across multiple languages.
LaBSE Language-agnostic BERT Sentence Embedding Feng et al. (2022) is trained with translation ranking loss and negative samples.
@@ 3.2 CLSD Results
The results of the models on the CLSD datasets are shown in Table 2. Regarding the direct cross-lingual evaluation, the Bitext Mining specialist LaBSE is the best model in this task with an average score of 94.43 (R@1), demonstrating its consistent ability to cross-lingually rank exact parallel text higher than the distractors. Another pattern that emerges is that models (M-E5-b/l, M-GTE) that were also trained on English-centric retrieval datasets, such as the MS Macro Nguyen et al. (2016), perform worse within our stricter task.
In the lower half of the table, we evaluate the models using English as a pivot language on both sides, employing the M2M 1.2B translation model Fan et al. (2021). Both LaBSE and M-MPNet show a performance decrease of 2.6/2.8 points compared to direct cross-lingual evaluation. In contrast, models with weaker performance in direct cross-lingual show only minor improvements, except the M-E5-b model, which gains 3.7 points.
@@ 3.3 Fine-grained similarity analysis
How do minor surface changes affect the semantic similarity between sentence pairs? Can linguistic analysis provide a more precise characterisation of the changes? To address these questions, we pick original-distractor pairs in which the distractor differs from the original by only a single word. This subset includes 523 pairs (5.5%) in German and 324 pairs (3.4%) in French. Using Stanza Qi et al. (2020) and further validation by a linguist annotator, we assign part-of-speech tags to the differing token. We measure the cross-lingual semantic change by subtracting the cosine distance between the original source sentence and the target distractor from the cosine distance between the original parallel pair. To improve comparability in our analysis across different models, we normalize222The average cosine similarity difference between parallel and unrelated cross-lingual sentence pairs in the original WMT19/21 datasets is the normalization factor in Figure 2. this score by dividing with for the corresponding model. Formally, we measure:

### v2

@@ Abstract
The evaluation of cross-lingual semantic search capabilities of models is often limited to existing datasets from tasks such as information retrieval and semantic textual similarity. To allow for domain-specific evaluation, we introduce Cross Lingual Semantic Discrimination (CLSD), a novel cross-lingual semantic search task that does not require a large evaluation corpus, only parallel sentences of the language pair of interest within the target domain. This task focuses on the ability of a model to cross-lingually rank the true parallel sentence higher than challenging distractors generated by a large language model. We create a case study of our introduced CLSD task for the language pair German-French in the news domain. Within this case study, we find that models that are also fine-tuned for retrieval tasks benefit from pivoting through English, while bitext mining models perform best directly cross-lingually. A fine-grained similarity analysis enabled by our distractor generation strategy indicate that different embedding models are sensitive to different types of perturbations.
@@ 1 Introduction
When considering which model to use to perform semantic search within a specific domain and language pair, standard benchmarks such as MTEB (Muennighoff et al., 2023) do not cover every domain within all language pairs. We posit that one should aim to directly evaluate semantic search for the target language pair within the domain of interest but the lack of dataset for such evaluation cases makes it difficult to take fundamental decisions for the application uses, such as the choice of an embedding model or whether to use English as a pivot language to perform the task.
@@ 2 Related Work
PAWS-X (Yang et al., 2019; Zhang et al., 2019) generated adversarial examples by rule-based and backtranslation methods to measure the limitations of paraphrase identification models. More recently, XSim++ (Chen et al., 2023) used rule-based text augmentation to generate synthetic examples for tuning LASER models (replacements by antonyms, swapping entities). Within semantic search, InPars (Bonifacio et al., 2022) use solely in-domain synthetic data to finetune retrieval models that surpass strong baselines.
@@
| Machine Translation (M2M 1.2B (Fan et al., 2021)) to EN -> CLSD Evaluation | ||||||||
@@ 3.1 Cross-Lingual Semantic Discrimination (CLSD)
In this paper, we test the hypothesis that LLMs can generate difficult examples that look similar on the surface, but convey different meanings. We prompt GPT-4 (gpt-4-0613) (OpenAI, 2023) to monolingually generate four distractors for each original sentence (see Figure 1). The distractors are requested to be structurally and lexically similar, whilst being semantically dissimilar. For each language, we manually examine 200 distractors of the original sentences and find that over 98% of the distractors meet these criteria. Example sentences are shown in Table 1.
@@
We generate and validate (samples of) four datasets, consisting of both language directions of German/French for the WMT19 (Barrault et al., 2019) and WMT21 (Akhbardeh et al., 2021) test sets.111The data is available at link. In the 18,928 generated distractors, we observe that the LLM used different modification strategies such as single polarity inversion, entity replacement or even multiple swaps with word-level Jaccard similarity between originals and distractors at (, ). Table 4 illustrates samples of distractors sets.
@@
To measure performance of embedding models, we use Precision@1 (P@1), defined as the proportion of samples in which the model ranks the true translation as more similar to the source sentence than the four distractors. For example, in Table 2, the column WMT19-DE->FR reports the P@1 of identifying the original French target sentence as more similar than its French distractors.
@@
We examine the following five embedding models available through SentenceTransformers (Reimers and Gurevych, 2019).
multilingual-e5-base/-large (M-E5-b/l) Wang et al. (2024, 2024) transform XLM-RoBERTa (Conneau et al., 2020) into a bi-encoder through weakly supervised contrastive pre-training with multilingual data and a second step of supervised fine-tuning on retrieval datasets.
@@
paraphrase-multilingual-mpnet-base-v2 (M-MPNet) Reimers and Gurevych (2020) use knowledge distillation through cosine loss training on parallel sentences between an English paraphrase trained teacher model (paraphrase-mpnet-base-v2) and a multilingual student model (XLM-RoBERTa (Conneau et al., 2020)) to teach the student to represent sentences closely across multiple languages.
LaBSE Language-agnostic BERT Sentence Embedding (Feng et al., 2022) is trained with translation ranking loss and negative samples.
@@ 3.2 CLSD Results
The results of the models on the CLSD datasets are shown in Table 2. Regarding the direct cross-lingual evaluation, the Bitext Mining specialist LaBSE is the best model in this task with an average score of 94.43 (P@1), demonstrating its consistent ability to cross-lingually rank exact parallel text higher than the distractors. Another pattern that emerges is that models (M-E5-b/l, M-GTE) that were also trained on English-centric retrieval datasets, such as the MS Macro (Nguyen et al., 2016), perform worse within our stricter task.
In the lower half of the table, we evaluate the models using English as a pivot language on both sides, employing the M2M 1.2B translation model (Fan et al., 2021). Both LaBSE and M-MPNet show a performance decrease of 2.6/2.8 points compared to direct cross-lingual evaluation. In contrast, models with weaker performance in direct cross-lingual show only minor improvements, except the M-E5-b model, which gains 3.7 points.
@@ 3.3 Fine-grained similarity analysis
How do minor surface changes affect the semantic similarity between sentence pairs? Can linguistic analysis provide a more precise characterisation of the changes? To address these questions, we pick original-distractor pairs in which the distractor differs from the original by only a single word. This subset includes 523 pairs (5.5%) in German and 324 pairs (3.4%) in French. Using Stanza (v 1.8) (Qi et al., 2020) and further validation by a linguist annotator, we assign part-of-speech tags to the differing token. We measure the cross-lingual semantic change by subtracting the cosine distance between the original source sentence and the target distractor from the cosine distance between the original parallel pair. To improve comparability in our analysis across different models, we normalize222The average cosine similarity difference between parallel and unrelated cross-lingual sentence pairs in the original WMT19/21 datasets is the normalization factor in Figure 2. this score by dividing with for the corresponding model. Formally, we measure: